PREFACE

This volume presents the proceedings of the international conference “Function Spaces, Differential Operators and Nonlinear Analysis” held in Paseky nad Jizerou, Krkono�e (Giant Mountains), Czech Republic, from 3 to 9 September 1995.

The proceedings contain 11 invited papers—including the paper by Olli Martio who at the very last moment had to cancel his participation at the conference—and 15 shorter communications. Almost all papers and communications fall within the three fields listed in the title and many of them reflect the tight connection between these fields.

The conference was already third in the series of meetings organized under the same name. The series was initiated to develop the existing contacts between mathematicians from Finland, German Democratic Republic and Czechoslovakia and to help to overcome the travelling obstructions. The first conference was held in Sodankylä in Finnish Lapland, in August 1988, the second one was organized in Friedrichroda, Thuringia, Germany, in September 1992. The proceedings were published in Pitman Research Notes in Mathematics Series, vol. 211, Longman Scientific & Technical, Harlow, 1989, (L. Päivärinta Ed.) and in Teubner Texte zur Mathematik, vol. 133, B. G. Teubner Verlagsgesellschaft, Stuttgart-Leipzig, 1993 (H.-J. Schmeisser and H. Triebel Eds.), respectively. The circle was closed with the conference in Paseky, the original travelling obstructions have in the meantime disappeared, and two of the three countries no longer exist. However, the conferences proved their raison d’etre and, hopefully, the next one will be held again in Finland.

We owe our thanks not only to the authors of the papers but also to all 43 participants who helped to create a friendly and creative atmosphere and contributed to the success of the meeting. The conference was organized by members of the Prague group around the seminar on theory of function spaces originated by Prof. Alois Kufner. It is a pleasure to express my gratitude to all of them for their support, in particular to Lubo� Pick and Bohum�r Opic who took on the main part of the organization. Special thanks are due to Blanka Koubov�, secretary of the organizing committee, who retyped and converted into the AMS-TEX most of the text, and to Karel Hor�k, who helped us to prepare the volume for printing.

Praha, April 1996 Ji�� R�kosn�k